OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the two smaller numbers is 22 less than the square of the larger number.

OpenStudy (anonymous):

1, 3 and 5

OpenStudy (anonymous):

let the numbers be a,a-2,a+2 then a*(a-2)+22=(a+2)^2 solcving this we get, 6a=18 nd a =3 so answer is 1,3,5